# Tracking Protostellar Variability in Massive Protoclusters with ALMA: I. Insights from QUARKS and MaMMOtH

### v1

@@ Abstract
Millimeter/Submillimeter variability is often attributed to dynamical disk-mediated accretion yet detection is limited to low-mass protostars in nearby clouds. Recent observations have also revealed significant (sub)millimeter variability in high-mass protostars, but the confirmed cases are scarce and lack systematic monitoring. In this work, we analyzed multi-epoch Atacama Large Millimeter/submillimeter Array (ALMA) Band 6 (1.3 mm) continuum observations of 22 massive protoclusters, with epoch separations ranging from a few hours to more than two years, while achieving a consistent angular resolution of ″. These data allow us to track variability of protostars across a broader mass range and in an environment markedly different from nearby clouds. Using a custom processing pipeline of data reduction, image alignment, and relative flux calibration, we achieve high-precision flux measurements and, for the first time, investigate millimeter variability in massive protoclusters by interferometric data from statistically investigate. Applying the astrodendro algorithm, we identified 383 condensations and tracked their variations in their peak intensities. Standard deviation analysis and difference maps reveal five variable sources, corresponding to a lower limit of 1.3% on the variable fraction. Among these, I13111–6228 stands out as it hosts a hypercompact Hii region that exhibits a 68% increase in continuum peak intensity over one year, with an uncertainty of 2%. This supports the burst mode of accretion picture in massive star formation as a viable route for the formation of massive stars.
@@ 1 Introduction
Understanding how protostars gain the bulk of their mass during the main accretion phase while they are still deeply embedded in their natal envelopes remains a fundamental challenge in star formation (e.g. 2009Evans; 2018Kristensen). Early estimates of protostellar accretion rates, inferred from luminosity under the assumption of steady accretion, were insufficient to explain the observed stellar masses within the expected formation timescales (approximately 0.1 Myr; 1990Kenyon; 2010Dunham). Although extended formation timescales and improved sensitivity have partially alleviated this “luminosity problem” (e.g., 2011Offner; 2023Fischer), the observed protostellar luminosities still vary by 3–4 orders of magnitude. This wide range has motivated a growing interest in time-variable accretion as a key mechanism for mass assembly.
@@
For low-mass young stellar objects, temporal photometric variability has been observed not only in the optical and infrared toward FU Orionis and EX Lupi-type objects in the optically revealed stage (2014Audard; 2025OYCAT; 2025Contreras), but also in the (sub)millimeter wavelengths at much earlier, embedded (Class 0, Class I) protostellar phases (2015Safron; 2024Mairs; 2025Sheehan; 2025Laznevoi). These outburst phenomena can be attributed to various gravitational instabilities in circumstellar disks, and, in particular, to gravitational instability in which the disk fragments into dense clumps that subsequently migrate inward and trigger episodic accretion onto the central protostar. Numerical simulations suggest such variable accretion events occurs not only in low-mass star formation (e.g., 2005Vorobyov; 2011Machida; 2012Nayakshin; 2015Vorobyov; 2016Hosokawa) but also in the high-mass star regime (e.g., 2017Meyer; 2018Meyer). Further studies suggest that massive protostars also undergo episodic accretion events that produce substantial luminosity bursts, and may contribute up to of their zero-age main-sequence mass, with burst durations ranging from several years to nearly a century (e.g., 2019Meyer; 2021Elbakyan).
@@
Massive protoclusters, which are forming compact clusters of protostars including high-mass ones with characteristic sizes of – pc (e.g., 2007Cyganowski; 2013Palau; 2014Monge; 2024Xu), are promising targets to study such variability in environments markedly different from nearby clouds. Their cluster members usually span a wide range of evolutionary stages, from ultracompact H ii regions and hot molecular cores to cold dust condensations (e.g., 2017ApJ...849...25L; 2021MNRAS.505.2801L; 2025ApJS..280...33Y), providing an ideal laboratory for investigating accretion-driven variability across different stages of massive star formation.
@@
Recent observations have directly confirmed powerful luminosity outbursts in massive star-forming regions, widely interpreted as episodes of enhanced accretion. The earliest confirmed examples are S255IR-NIRS3 (2017Caratti; 2018Cesaroni; 2018Szymczak; 2019Liu; 2020Liu; 2020Uchiyama) and NGC 6334I-MM1 (2017Hunter; 2018Hunter; 2021Hunter; 2018Brogan; 2018MacLeod; 2021Chibueze). Multi-wavelength observations showed that the continuum emission from millimeter to infrared wavelengths increased by a factor of 5.5 and 16.3, respectively. These events are interpreted as episodic accretion bursts of the internal massive young stellar objects (MYSOs), which are likely triggered by disk fragmentation, followed by inward migration of the clumps and their tidal destruction. (e.g., 2017Meyer; 2021Elbakyan). Such behavior is consistent with the theoretical burst mode of accretion proposed for massive star formation (2021Meyer; 2022Meyer).
@@
Notably, both the S255IR-NIRS3 and NGC 6334I-MM1 outbursts were accompanied by flaring of the 6.7 GHz methanol (CH3OH) maser (2015Fujisawa; 2018Szymczak; 2018MacLeod; 2019Brogan). Since Class II CH3OH masers, including the 6.7 GHz transition, are radiatively pumped by far-infrared photons (2003Minier; 2008Xu), the observed maser flaring indicates an increase in the far-infrared thermal radiation field within the envelope surrounding the MYSO, implying enhanced heating of the circumstellar material. Adding to this evidence are confirmed cases like G358.930.03 MM1 (2021Stecklum; 2020Burns; 2023Burns), G24.33+0.14 (2022Hirota), G323.460.08 (2024Wolf), and M17 MIR (2021Chen; 2024Zhou; 2025Chen).
@@
These findings further support the idea that episodic accretion may play a crucial role in the formation of massive stars: it governs their early formation phase, determines their stellar structure at the zero-age main-sequence stellar their strong line-driven winds start to be expelled from the stellar surface, and eventually influences their entire stellar evolution, core-collapse characteristics, and resulting pulsar properties. Despite their increasing number, accretion bursts in massive protostars remain rare and largely serendipitous, hindering efforts to statistically constrain the nature of episodic accretion in the high-mass regime. Therefore, systematic, long-term monitoring is essential to understand their occurrence rate, properties, and driving mechanisms.
@@
The JCMT Transient Survey (2017Herczeg; 2018Johnstone; 2024Mairs) observed six fields targeting at intermediate- to high-mass star-forming regions. However, the relatively large beam size imposes limitations on detailed variability monitoring in such distant regions (2025Chen; 2024Park; Wang_prep; Zhang_prep). Previous systematic interferometric studies have investigated (sub)millimeter variability in nearby low-mass star-forming regions using facilities such as the Submillimeter Array (SMA), the Combined Array for Research in Millimeter-wave Astronomy (CARMA), the NOrthern Extended Millimeter Array (NOEMA) and ALMA (e.g., 2018Liu; 2019Francis; 2022Francis; 2020Wendeborn; 2023Vargas), demonstrating the feasibility of interferometric monitoring of variability. Yet such efforts remain confined to nearby molecular clouds, and a systematic interferometric search for variability across multiple massive protoclusters has not yet been undertaken.
@@
In this paper, we present the first systematic investigation of continuum variability in massive protoclusters at 1.3 mm wavelength, based on high-resolution interferometric data obtained from two recent ALMA projects: the Querying Underlying mechanisms of massive star formation with ALMA-Resolved gas Kinematics and Structures (QUARKS) survey (Liu2024QUARKS-I) and the Massive Star-Forming Regions with Variable Methanol Masers: Observations at High Angular Resolution (MaMMOtH) survey (Liu et al. in prep.). Our analysis includes multi-epoch observations of 22 distant massive protoclusters. Each source has observations from at least two epochs separated by more than one year, while across the full dataset the epoch separations range from only a few hours to more than two years. Specifically, this work expands upon previous studies of (sub)millimeter continuum variability by searching for variable sources across a wider range of masses within more massive, active, and complex star-forming environments.
@@
The paper is structured as follows. In Sect. 2, we describe source selection and present the sample. We also provide details of the corresponding ALMA observations. In Sect. 3, we outline each step of the data reduction process for the multi-epoch continuum maps. This includes calibration, identification of line-free channels for continuum determination, self-calibration, and smoothing all maps to a common synthesized beam to ensure consistent angular resolution across epochs. In Sect. 4, we present the procedures for source identification and flux extraction across multiple epochs. We also describe the relative flux calibration methods applied to ensure consistency, followed by an analysis of the resulting millimeter variability. In Sect. 5, we discuss the results, outline the current limitations of this study, and present our plans for future investigations. Finally, in Sect. 6, we summarize our main findings and conclusions.
2 Observations and Sample Selections
@@
A schematic overview of the full processing workflow is provided in Figure 1. This study uses data from two independent ALMA projects: QUARKS and MaMMOtH. The QUARKS Survey (Project IDs: 2019.1.00685.S and 2021.1.00095.S; PIs: Lei Zhu, Guido Garay, and Tie Liu) investigates 139 massive star-forming clumps across 156 pointings using ALMA Band 6 observations (). The QUARKS observations began in late October 2021 using both the ALMA 12-m array configurations (C-2 and C-5) and the Atacama Compact Array (ACA) 7-m antennas. The ACA observations were completed in late May 2022 (2024Xu), followed by C-2 and C-5 executions that continued through June 2024. The temporal coverage and per-epoch sensitivity of the QUARKS observations are sufficient to enable internal variability comparisons within the survey. As a follow-up to the ALMA Three-millimeter Observations of Massive Star-forming regions survey (ATOMS; Liu2020ATOMS-I), QUARKS aims to statistically characterize key processes of star formation (e.g., fragmentation, outflows, disks) within an unbiased sample of protoclusters. Detailed descriptions of survey design, target selection, and observation strategy are provided in Liu2024QUARKS-I.
@@
The MaMMOtH Survey (Project IDs: 2021.1.00311.S and 2022.1.00974.S; PI: Sheng-Yuan Liu) targets 169 massive star-forming regions associated with Class II CH3OH masers. The sample is primarily drawn from two well-established catalogs of monitoring observations of 6.7 GHz methanol masers (2004Goedhart; 2018Szymczak), and is supplemented by two additional well-known maser sources, G352.6301.067 (2019Chen) and G353.2730.641 (2013Motogi), based on previous detections of CH3OH maser activity. Observations were conducted using ALMA Band 6 with the 12-m array configurations (C-4 and C-5), complemented by Band 7 observations with the ACA. The ACA observations began in October 2021 and were completed in mid-May 2023, while the C-4 executions were carried out from April 2022 to April 2024, followed by the C-5 observations that continued from May 2022 through June 2024. The primary goal of the MaMMOtH Survey is to establish a statistical baseline for the study of millimeter continuum variability in massive star-forming regions associated with 6.7 GHz methanol masers, and to examine its connection to the physical and chemical conditions of the star-forming environment. The survey description, including full observational details, will be introduced in Liu et al. (in prep.).
@@
The ALMA Band 6 receivers in dual-polarization mode were used to conduct both the QUARKS and MaMMOtH surveys. For the QUARKS survey, four spectral windows were configured with bandwidths of 1.875 GHz and a velocity resolution of per spectral window (SPW). The central frequencies of the spectral windows were set at approximately 217.92, 220.32, 231.37, and 233.52 GHz. Detailed information on the spectral setup and targeted molecular lines can be found in Table 2 of Liu2024QUARKS-I. For the MaMMOtH survey, four spectral windows were similarly configured, each with a bandwidth of 1.875 GHz and a velocity resolution of per SPW. The central frequencies of the spectral windows were set at approximately 217.63, 220.00, 231.05, and 232.87 GHz.
@@
Benefiting from the comparable frequency coverage and matched observational setups (ALMA C-5 configuration and resolution) of the two projects, we constructed a multi-epoch sample of 22 massive protoclusters from the QUARKS and MaMMOtH surveys. We cross-matched the MaMMOtH and QUARKS source lists by requiring that their pointing centers differ by less than 1″, and this yielded 22 massive protoclusters with overlapping coverage. Among these cross-matched sources, eight massive protoclusters have at least two epochs from the two surveys that are separated by more than one year. In addition, the QUARKS survey alone provides 14 sources with two epochs separated by more than one year. Taken together, these 22 massive protoclusters constitute the largest sample to date of sources with observational epochs separated by more than one year across the two projects. Additional shorter-interval epochs are included when the observational setups were consistent. The molecular gas reservoirs of these protoclusters range from 68 to 7585 , and they are located at distances of 1.4 to 11.6 kpc (Liu2024QUARKS-I). A summary of the observational parameters, including observing dates, calibrators, and baseline ranges, is provided in Appendix A.
3 Data Reduction Methods
3.1 Spectral Line Flagging and Self-calibration

### v3

@@ Abstract
Millimeter/Submillimeter variability is often attributed to dynamical disk-mediated accretion yet detection is limited to low-mass protostars in nearby clouds. Recent observations have also revealed significant (sub)millimeter variability in high-mass protostars, but the confirmed cases are scarce and lack systematic monitoring. In this work, we analyzed multi-epoch Atacama Large Millimeter/submillimeter Array (ALMA) Band 6 (1.3 mm) continuum observations of 22 massive protoclusters, with epoch separations ranging from a few hours to more than two years, while achieving a consistent angular resolution of ″. These data allow us to track variability of protostars across a broader mass range and in an environment markedly different from nearby clouds. Using a custom processing pipeline of data reduction, image alignment, and relative flux calibration, we achieve high-precision flux measurements and, for the first time, investigate millimeter variability in massive protoclusters using interferometric data from the statistical point of view. Applying the astrodendro algorithm, we identified 383 condensations and tracked their variations in their peak intensities. Standard deviation analysis and difference maps reveal five variable sources, corresponding to a lower limit of 1.3% on the variable fraction. Among these, I13111–6228 stands out as it hosts a hypercompact Hii region that an increase of approximately 68% in continuum peak intensity over one year, with an uncertainty of 2%.
@@ I Introduction
Understanding how protostars gain the bulk of their mass during the main accretion phase while they are still deeply embedded in their natal envelopes remains a fundamental challenge in star formation (e.g. Evans et al. 2009; Kristensen and Dunham 2018). Early estimates of protostellar accretion rates, inferred from luminosity under the assumption of steady accretion, were insufficient to explain the observed stellar masses within the expected formation timescales (approximately 0.1 Myr; Kenyon et al. 1990; Dunham et al. 2010). Although extended formation timescales and improved sensitivity have partially alleviated this “luminosity problem” (e.g., Offner and McKee 2011; Fischer et al. 2023), the observed protostellar luminosities still vary by 3–4 orders of magnitude. This wide range has motivated a growing interest in time-variable accretion as a key mechanism for mass assembly.
@@
For low-mass young stellar objects, temporal photometric variability has been observed not only in the optical and infrared toward FU Orionis and EX Lupi-type objects in the optically revealed stage (Audard et al., 2014; Contreras Peña et al., 2025a, b), but also in the (sub)millimeter wavelengths at much earlier, embedded (Class 0, Class I) protostellar phases (Safron et al., 2015; Mairs et al., 2024; Sheehan et al., 2025; Laznevoi et al., 2025). These outburst phenomena can be attributed to various gravitational instabilities in circumstellar disks, and, in particular, to gravitational instability in which the disk fragments into dense clumps that subsequently migrate inward and trigger episodic accretion onto the central protostar. Numerical simulations suggest such variable accretion events occurs not only in low-mass star formation (e.g., Vorobyov and Basu 2005; Machida et al. 2011; Nayakshin and Lodato 2012; Vorobyov and Basu 2015; Hosokawa et al. 2016) but also in the high-mass star regime (e.g., Meyer et al. 2017, 2018). Further studies suggest that massive protostars also undergo episodic accretion events that produce substantial luminosity bursts, and may contribute up to of their zero-age main-sequence mass, with burst durations ranging from several years to nearly a century (e.g., Meyer et al., 2019; Elbakyan et al., 2021).
@@
Massive protoclusters, which are forming compact clusters of protostars including high-mass ones with characteristic sizes of – pc (e.g., Cyganowski et al., 2007; Palau et al., 2013; Sánchez-Monge et al., 2014; Xu et al., 2024), are promising targets to study such variability in environments markedly different from nearby clouds. Their members typically span a wide range of evolutionary stages, from starless cores through hot molecular cores to ultracompact Hii regions (e.g., Liu et al., 2017, 2021; Yang et al., 2025), providing an ideal laboratory for investigating accretion-driven variability across different stages of massive star formation.
@@
Recent observations have directly confirmed powerful luminosity outbursts in massive star-forming regions, widely interpreted as episodes of enhanced accretion. The earliest confirmed examples are S255IR-NIRS3 (Caratti o Garatti et al., 2017; Cesaroni et al., 2018; Szymczak et al., 2018; Liu et al., 2019, 2020a; Uchiyama et al., 2020) and NGC 6334I-MM1 (Hunter et al., 2017, 2018, 2021; Brogan et al., 2018; MacLeod et al., 2018; Chibueze et al., 2021). Multi-wavelength observations showed that the continuum emission from millimeter to infrared wavelengths increased by a factor of 5.5 and 16.3. These events are interpreted as episodic accretion bursts of the internal massive young stellar objects (MYSOs), which are likely triggered by disk fragmentation, followed by inward migration of the clumps and their tidal destruction. (e.g., Meyer et al., 2017; Elbakyan et al., 2021). Such behavior is consistent with the theoretical burst mode of accretion proposed for massive star formation (Meyer et al., 2021, 2022).
@@
Notably, both the S255IR-NIRS3 and NGC 6334I-MM1 outbursts were accompanied by flaring of the 6.7 GHz methanol (CH3OH) maser (Fujisawa et al., 2015; Szymczak et al., 2018; MacLeod et al., 2018; Brogan et al., 2019). Since Class II CH3OH masers, including the 6.7 GHz transition, are radiatively pumped by far-infrared photons (Minier et al., 2003; Xu et al., 2008), the observed maser flaring indicates an increase in the far-infrared thermal radiation field within the envelope surrounding the MYSO, implying enhanced heating of the circumstellar material. Adding to this evidence are confirmed cases like G358.93-0.03 MM1 (Stecklum et al., 2021; Burns et al., 2020, 2023), G24.33+0.14 (Hirota et al., 2022), G323.46-0.08 (Wolf et al., 2024), and M17 MIR (Chen et al., 2021; Zhou et al., 2024; Chen et al., 2025).
@@
These findings further support the idea that episodic accretion may play a crucial role in the formation of massive stars. By assembling a substantial fraction of stellar mass in short-lived accretion bursts, this process modulates radiative feedback and the thermal and ionization structure of the circumstellar environment, producing the large luminosity spread and characteristic multi-wavelength variability observed in massive protostars. Despite their increasing number, accretion bursts in massive protostars remain rare and largely serendipitous, hindering efforts to statistically constrain the nature of episodic accretion in the high-mass regime. Therefore, systematic, long-term monitoring is essential to understand their occurrence rate, properties, and driving mechanisms.
@@
The JCMT Transient Survey (Herczeg et al., 2017; Johnstone et al., 2018; Mairs et al., 2024) observed six fields targeting at intermediate- to high-mass star-forming regions. However, the relatively large beam size imposes limitations on detailed variability monitoring in such distant regions (Chen et al., 2025; Park et al., 2024; Wang and others, in prep.; Zhang and others, in prep.). Previous systematic interferometric studies have investigated (sub)millimeter variability in nearby low-mass star-forming regions using facilities such as the Submillimeter Array (SMA), the Combined Array for Research in Millimeter-wave Astronomy (CARMA), the NOrthern Extended Millimeter Array (NOEMA) and ALMA (e.g., Liu et al., 2018; Francis et al., 2019, 2022; Wendeborn et al., 2020; Vargas-González et al., 2023), demonstrating the feasibility of interferometric monitoring of variability. Yet such efforts remain confined to nearby molecular clouds, and a systematic interferometric search for variability across multiple massive protoclusters has not yet been undertaken.
@@
In this paper, we present the first systematic investigation of continuum variability in massive protoclusters at 1.3 mm wavelength, based on high-resolution interferometric data obtained from two recent ALMA projects: the Querying Underlying mechanisms of massive star formation with ALMA-Resolved gas Kinematics and Structures (QUARKS) survey (Liu et al., 2024) and the Massive Star-Forming Regions with Variable Methanol Masers: Observations at High Angular Resolution (MaMMOtH) survey (Liu and others, in prep.). Our analysis includes multi-epoch observations of 22 distant massive protoclusters. Each source has observations from at least two epochs separated by more than one year, while across the full dataset the epoch separations range from only a few hours to more than two years. Specifically, this work expands upon previous studies of (sub)millimeter continuum variability by searching for variable sources across a wider range of masses within more massive, active, and complex star-forming environments.
@@
The paper is structured as follows. In Sect. II, we describe source selection and present the sample. We also provide details of the corresponding ALMA observations. In Sect. III, we outline each step of the data reduction process for the multi-epoch continuum maps. This includes calibration, identification of line-free channels for continuum determination, self-calibration, and smoothing all maps to a common synthesized beam to ensure consistent angular resolution across epochs. In Sect. IV, we present the procedures for source identification and flux extraction across multiple epochs. We also describe the relative flux calibration methods applied to ensure consistency, followed by an analysis of the resulting millimeter variability. In Sect. V, we discuss the results, outline the current limitations of this study, and present our plans for future investigations. Finally, in Sect. VI, we summarize our main findings and conclusions.
II Observations and Sample Selections
@@
A schematic overview of the full processing workflow is provided in Figure 1. This study uses data from two independent ALMA projects: QUARKS and MaMMOtH. The QUARKS Survey (Project IDs: 2019.1.00685.S and 2021.1.00095.S; PIs: Lei Zhu, Guido Garay, and Tie Liu) investigates 139 massive star-forming clumps that have IRAS colors similar to those of ultracompact H ii regions, using ALMA Band 6 observations ( mm) across 156 pointings. The QUARKS observations began in late October 2021 using both the ALMA 12-m array configurations (C-2 and C-5) and the Atacama Compact Array (ACA) 7-m antennas. The ACA observations were completed in late May 2022 (Xu et al., 2024), followed by C-2 and C-5 executions that continued through June 2024. The temporal coverage and per-epoch sensitivity of the QUARKS observations are sufficient to enable internal variability comparisons within the survey. As a follow-up to the ALMA Three-millimeter Observations of Massive Star-forming regions survey (ATOMS; Liu et al., 2020b), QUARKS aims to statistically characterize key processes of star formation (e.g., fragmentation, outflows, disks) within an unbiased sample of protoclusters. Detailed descriptions of survey design, target selection, and observation strategy are provided in Liu et al. (2024).
@@
The MaMMOtH Survey (Project IDs: 2021.1.00311.S and 2022.1.00974.S; PI: Sheng-Yuan Liu) targets 169 massive star-forming regions associated with Class II CH3OH masers. The sample is primarily drawn from two well-established catalogs of monitoring observations of 6.7 GHz methanol masers (Goedhart et al., 2004; Szymczak et al., 2018), and is supplemented by two additional well-known maser sources, G352.630-1.067 (Chen et al., 2019) and G353.2730.641 (Motogi et al., 2013), based on previous detections of Class II CH3OH maser activity. Observations were conducted using ALMA Band 6 with the 12-m array configurations (C-4 and C-5), complemented by Band 7 observations with the ACA. The ACA observations began in October 2021 and were completed in mid-May 2023, while the C-4 executions were carried out from April 2022 to April 2024, followed by the C-5 observations that continued from May 2022 through June 2024. The primary goal of the MaMMOtH Survey is to establish a statistical baseline for the study of millimeter continuum variability in massive star-forming regions associated with 6.7 GHz methanol masers, and to examine its connection to the physical and chemical conditions of the star-forming environment. The survey description, including full observational details, will be introduced in Liu et al. (in prep.).
@@
The ALMA Band 6 receivers in dual-polarization mode were used to conduct both the QUARKS and MaMMOtH surveys. For the QUARKS survey, four spectral windows (SPWs) were configured, each with a bandwidth of 1.875 GHz and a velocity resolution of . The central frequencies of the SPWs were set at approximately 217.92, 220.32, 231.37, and 233.52 GHz. Detailed information on the spectral setup and targeted molecular lines can be found in Table 2 of Liu et al. (2024). For the MaMMOtH survey, four spectral windows were similarly configured, each with a bandwidth of 1.875 GHz and a velocity resolution of . The central frequencies of the SPWs were set at approximately 217.63, 220.00, 231.05, and 232.87 GHz.
@@
Benefiting from the comparable frequency coverage and matched observational setups (ALMA C-5 and resolution) of the two projects, we constructed a multi-epoch sample of 22 massive protoclusters from the QUARKS and MaMMOtH surveys. We cross-matched the MaMMOtH and QUARKS source lists by requiring that their pointing centers differ by less than 60″, and this yielded 22 massive protoclusters with overlapping coverage. Among these cross-matched sources, eight massive protoclusters have at least two epochs from the two surveys that are separated by more than one year. In addition, the QUARKS survey alone provides 14 sources with two epochs separated by more than one year. Taken together, these 22 massive protoclusters constitute the largest sample to date of sources with observational epochs separated by more than one year across the two projects. Additional shorter-interval epochs are included when the observational setups were consistent. The molecular gas reservoirs of these protoclusters range from 68 to 7585 , and they are located at distances of 1.4 to 11.6 kpc (Liu et al., 2024; Xu et al., 2024). A summary of the observational parameters, including observing dates, calibrators, and baseline ranges, is provided in Appendix A.
III Data Reduction Methods
III.1 Spectral Line Flagging and Self-calibration